# Enhanced Visual Question Answering: A Comparative Analysis and Textual Feature Extraction Via Convolutions

### v1

@@ 4.2 Experimental Settings
  • 2Layer-LSTM, BiGRU, BiLSTM: Extended from the One-Layer GRU and One-layer LSTM by simply adding bidirectional processing or an extra layer.
@@
ConvGRU Settings. For our ConvGRU model (Section 3.4), we experimented with different kernel sizes (1, 2, 3, and 4) for convolutional layers to see how they affect text feature extraction. Following the idea from [31], we used multiple kernel sizes to capture text features at various scales. Instead of pooling, we chose to concatenate features from different kernels to avoid losing too much textual information since most input questions are quite short. Our main focus was on multiscale kernels, specifically combinations of k=2+3 and k=1+2+3, to leverage multiple text feature representations.
@@ 4.3 Results
Table 2: Comparison of different textual models. ’SA’ and ’MHA’ refer to Self-Attention and Multi-Head Self-Attention respectively, and all configurations are the same except for textual encoders.
@@
RCNN+2Layer-LSTM 48.23 55.74 64.98 22.34 25.81 36.76
@@
Table 3: Performance comparison of ConvGRU configurations with varying kernel sizes (denoted by numbers in parentheses) and the residual connections (indicated by ’Res’). The table presents VQA accuracy and Balanced Pair Accuracy across different question types (Yes/No, Number, Count, Other) and a cumulative score for all questions. ’Head’ and ’Tail’ refer to asymmetric head padding and asymmetric tail padding discussed in Section 3.4, respectively.
@@
On the contrary, local text features seem to play a more pivotal role in determining model accuracy considering the nature of questions in VQA-v2 dataset. This can be seen from the results shown in Table 3. The convolutional layer and residual unit in ConvGRUs tend to capture extra local dependencies within the text data effectively without losing information from initial text representations. When comparing the ConvGRU variants with kernel sizes ranging from 1 to 4 (Conv(1) to Conv(4)), a clear pattern emerges: The Conv(1) model, equally performing a linear projection, offers limited benefits in terms of capturing the nuanced relationships between words in a question. On the other hand, models with larger kernel sizes, like Conv(4), demonstrate a decline in performance, suggesting that excessively large kernels may overlook crucial local textual features. This observation implies that kernels larger than 4 are too broad, potentially diluting the model’s focus on the immediate contextual relevance between adjacent words. Conversely, kernel sizes of 2 and 3 strike a balance.
@@
Although the Conv(3+Res) configuration shows competitive, and in some aspects, superior performance to Conv(2+3+Res,Head), the latter’s broader applicability across different text scenarios, as further discussed in Section 4.6, positions it as a more versatile model. The Conv(2+3+Res,Head) model not only capitalizes on multi-scale convolution benefits but also enhances the interpretability and adaptability of the model to various question types within the VQA framework. This makes it an optimal choice for addressing the diverse challenges presented by the VQA tasks.
Also, it needs to point out that relatively incremental improvement by ConvGRU over the baseline GRU model may stem from limitations inherent in the Image Encoder, attention mechanisms, and the visual-textual fusion methods employed. So advancements in these areas might further unlock the potential of ConvGRU in VQA tasks.
4.4 Unexpected Shortfalls: Why Transformers Struggle with VQA Text
The underperformance of Transformer-based textual models in Visual Question Answering (VQA) tasks may initially appear surprising, given their success in broader NLP applications. This discrepancy arises from the unique textual characteristics within VQA, where questions are typically short and semantically similar. Unlike many NLP tasks where global contextual insights are crucial, VQA demands precise extraction of information from concise queries.
Transformer models excel in capturing relationships in extensive text sequences through their self-attention mechanisms, adept at uncovering deep, long-range dependencies. However, VQA tasks often involve brief textual prompts seeking specific visual details. This context significantly limits the utility of global contextual analysis, as the essence of VQA questions lies in localized cues rather than extensive narrative contexts. Furthermore, the semantic similarity among VQA questions adds another layer of complexity. Transformers, designed for distinguishing and leveraging nuanced differences across large texts, may not effectively prioritize the subtle distinctions critical in VQA. Thus, their architectural advantages, while transformative for general NLP, may inadvertently obscure the localized, specific cues vital for accurate Visual Question Answering.
In essence, the architectural design of Transformer models, though revolutionary for capturing global textual relationships, aligns less effectively with the requirements of VQA tasks. This misalignment underscores the necessity for approaches that emphasize local feature extraction, suggesting a tailored adaptation of text models to better accommodate the succinct and semantically concentrated nature of VQA queries.
@@ 4.5 How Short Can VQA Questions Get?
Figure 3: Distribution of question lengths within the VQA-v2 dataset. Red bars represent the total number of questions at each length, while blue bars show the number of counting questions that start with ’How many’. The majority of questions are observed to be under 10 words in length, with a significant portion of counting questions falling within the 5-8 word range. This underlines the dominance of short questions in the VQA-v2 dataset, corroborating the efficacy of using small kernels in convolutional approaches for extracting local text features.
We conduct an analysis of the length distribution of questions in the VQA-v2 dataset to demonstrate how the convolutional extraction of local features can lead to performance improvements. As shown in Figure 3, we use bar chart to represent the distribution, with the X-axis indicating the length of questions and the Y-axis indicating the frequency of questions at each length. In our comprehensive analysis of 658,111 questions from VQA-v2 dataset, we found that a significant majority, 636,894 questions (96.78%), have lengths ranging from 3 to 10 words. Notably, 569,437 questions (86.53%) are within the 4-8 word range. A steep decrease in question frequency is observed with increasing length. Focusing on the "How many" counting questions (72,012 in total), a predominant 61,516 questions (85.42%) fall within the 5-8 word range, with questions exceeding 10 words represent a mere 4.1% of counting queries. These insights highlight the predominance of short questions in VQA-v2 dataset, demonstrating the effectiveness of smaller convolution kernels for local feature extraction. The prevalence of question-type determinants at the start of queries ("How many", "Is there", "Do", etc.) further advocates for the superiority of compact kernels. This observation also aligns with the performance dip for k=4 scenarios depicted in Table 2, reinforcing smaller kernels’ utility in enhancing textual modality within VQA frameworks.
@@ 4.6 Qualitative Results
In this section, we choose four distinct cases from validation set to demonstrate how varying configurations of GRUs impact prediction accuracy (Figure 4). Our qualitative analysis underscores that the GRU model augmented with 2-gram and 3-gram features outperforms both the standard one-layer GRU and the version with only 3-gram features.
  • As shown in Case (a), the standard GRU incorrectly predicts the answer to be "blue", likely confusing the color of the background with that of the text. Conversely, both the ConvGRU(3) and the ConvGRU(2+3) models accurately identify the correct answer as "white". This illustrates the enhanced capability of these models to capture key phrase pairs such as "color of" and "wording on" through 2-grams, and rich context through 3-grams like "the color of" and "of the wording," effectively utilizing the convolutional layers to discern the finer details of the question.
  • In Case (b), the GRU model gives a completely wrong prediction of "picture", demonstrating a reliance on simplistic word-level features without leveraging n-gram characteristics. This limitation likely resulted in the model’s inability to grasp the relational and structural features of the sentence. On the other hand, the ConvGRU(3) model inferred "sitting"—a rationally deduced yet contextually incorrect response, recognizing 3-gram features that suggest a person’s state or action in a specific location. The ConvGRU(2+3), by incorporating both 2-gram and 3-gram features, correctly comprehends the query’s intent and the subject’s context, leading to an accurate "reading" prediction. This case exemplifies how combining different n-grams can enrich linguistic feature understanding and enhance alignment with image features for more accurate predictions.
  • In both the counting task and the yes/no task, as shown in Case (c) and Case (d), ConvGRU(2+3) showcases better performance. The ability to understand relational context and specific details allows it to correctly count "3" boys on the field and accurately confirm the presence of avocado on the sandwich with a "yes." These outcomes illustrate the effectiveness of integrating multi-gram features in enhancing the GRU’s comprehension.
Across all cases, the GRU model enhanced with both 2-gram and 3-gram features consistently yields the most accurate predictions, effectively capturing both the essential keyword pairs and the contextual information necessary for understanding the questions.
@@ 5 Conclusion
This paper explored the effectiveness of complex sequential models versus simpler models for capturing textual features in Visual Question Answering (VQA) tasks. Through our experiments, we found that the ConvGRU model, designed to capture local text features, slightly outperformed the baseline GRU model, while models employing advanced mechanisms such as Self-Attention and Transformer Encoders experienced a notable decrease in performance. While our findings offer valuable insights, it’s important to acknowledge that our study was not without its constraints. Specifically, we did not explore alternative methods for other components within the VQA framework, such as more advanced models for image feature extraction beyond the Residual Convolutional Neural Network (R-CNN) utilized in this study. This limitation suggests that there could be untapped potential in integrating more effective techniques that may further enhance the overall model performance in VQA tasks.
Looking forward, future research should focus on targeted improvements in both individual components and fusion methods from a multi-modal understanding perspective. This could involve the adoption of advanced neural architectures that are specifically designed for more effective handling of visual data, or the exploration of sophisticated fusion techniques that can seamlessly integrate visual and textual modalities, ensuring a more coherent and contextually relevant understanding in VQA tasks.

### v2

@@ 4.2.2 Compared Methods
  • 2-Layer-LSTM, BiGRU, BiLSTM: Extended from the One-Layer GRU and One-layer LSTM by simply adding bidirectional processing or an extra layer.
@@
4.2.3 ConvGRU Setting
For our ConvGRU model (Section 3.4), we experimented with different kernel sizes (1, 2, 3, and 4) for convolutional layers to see how they affect text feature extraction. Following the idea from [27], we used multiple kernel sizes to capture text features at various scales. Instead of pooling, we chose to concatenate features from different kernels to avoid losing too much textual information since most input questions are quite short. Our main focus was on multi-scale kernels, specifically combinations of k=2+3 and k=1+2+3, to leverage multiple text feature representations.
@@ 4.3 Results
Table 2: Comparison of different textual models. "SA" and "MHA" refer to Self-Attention and Multi-Head Self-Attention respectively, and all configurations are the same except for textual encoders.
@@
RCNN+2-Layer-LSTM 48.23 55.74 64.98 22.34 25.81 36.76
@@
Table 3: Performance comparison of ConvGRU configurations with varying kernel sizes (denoted by numbers in parentheses) and the residual connections (indicated by ’Res’). “Head” and “Tail” refer to asymmetric head padding and asymmetric tail padding discussed in Section 3.4, respectively.
@@
On the contrary, local text features seem to play a more pivotal role in determining model accuracy considering the nature of questions in VQA-v2 dataset. This can be seen from the results shown in Table 3. The convolutional layer and residual unit in ConvGRU tend to capture extra local dependencies within the text data effectively without losing information from initial text representations. When comparing the ConvGRU variants with kernel sizes ranging from 1 to 4 (Conv(1) to Conv(4)), a clear pattern emerges: The Conv(1) model, equally performing a linear projection, offers limited benefits in terms of capturing the nuanced relationships between words in a question. On the other hand, models with larger kernel sizes, like Conv(4), demonstrate a decline in performance, suggesting that excessively large kernels may overlook crucial local textual features. This observation implies that kernels larger than 4 are too broad, potentially diluting the model’s focus on the immediate contextual relevance between adjacent words. Conversely, kernel sizes of 2 and 3 strike a balance.
@@
Although the Conv(3+Res) shows competitive, and in some aspects, superior performance to Conv(2+3+Res,Head), the latter’s broader applicability across different text scenarios, as further discussed in Section 4.6, positions it as a more versatile model. The Conv(2+3+Res,Head) model not only capitalizes on multi-scale convolution benefits but also enhances the interpretability and adaptability of the model to various question types within the VQA framework. This makes it an optimal choice for addressing the diverse challenges presented by the VQA tasks.
Also, it needs to point out that relatively incremental improvement by ConvGRU over the baseline GRU model may stem from limitations inherent in the Image Encoder, attention mechanisms, and the visual-textual fusion methods employed. Therefore, advancements in these areas might further unlock the potential of ConvGRU in VQA tasks.
4.4 Unexpected Shortfalls: Why Transformer Encoder Struggles with VQA Text
The underperformance of Transformer Encoder in Visual Question Answering (VQA) tasks may initially appear surprising, given their success in broader NLP applications. This discrepancy arises from the unique textual characteristics within VQA, where questions are typically short and semantically similar. Unlike many NLP tasks where global contextual insights are crucial, VQA demands precise extraction of information from concise queries.
Transformer Encoder is good at capturing relationships in extensive text sequences through self-attention mechanisms. However, VQA tasks often involve brief textual prompts seeking specific visual details. This context significantly limits the utility of global contextual analysis, as the essence of VQA questions lies in localized cues rather than extensive narrative contexts. Furthermore, the semantic similarity among VQA questions adds another layer of complexity. Transformer Encoder may not effectively prioritize the subtle distinctions critical in VQA. Thus, their architectural advantages, while transformative for general NLP, may inadvertently obscure the localized, specific cues vital for accurate Visual Question Answering.
In essence, the architectural design of Transformer Encoder, though revolutionary for capturing global textual relationships, aligns less effectively with the requirements of VQA tasks. This misalignment underscores the necessity for approaches that emphasize local feature extraction, suggesting a tailored adaptation of text models to better accommodate the concise and semantically concentrated nature of VQA queries.
@@ 4.5 How Short Can VQA Questions Get?
Figure 3: Distribution of question lengths within the VQA-v2 dataset. Red bars represent the total number of questions at each length, while blue bars show the number of counting questions that start with ’How many’.
We conduct an analysis of the length distribution of questions in the VQA-v2 dataset to demonstrate how the convolutional extraction of local features can lead to performance improvements.
As shown in Figure 3, we use bar chart to represent the distribution, with the X-axis indicating the length of questions and the Y-axis indicating the frequency of questions at each length. In our comprehensive analysis of 658,111 questions from VQA-v2 dataset, we found that a significant majority, 636,894 questions (96.78%), have lengths ranging from 3 to 10 words. Notably, 569,437 questions (86.53%) are within the 4 to 8-word range. A steep decrease in question frequency is observed with increasing length. Focusing on the "How many" counting questions (72,012 in total), a predominant 61,516 questions (85.42%) fall within the 5 to 8-word range, with questions exceeding 10 words represent a mere 4.1% of counting queries.
These insights highlight the predominance of short questions in VQA-v2 dataset, demonstrating the effectiveness of smaller convolution kernels for local feature extraction. The prevalence of question-type determinants at the start of queries ("How many", "Is there", "Do", etc.) further advocates for the superiority of compact kernels. This observation also aligns with the performance dip for k=4 scenarios depicted in Table 3, reinforcing smaller kernels’ utility in enhancing textual modality within VQA frameworks.
@@ 4.6 Qualitative Results
In this section, we present four distinct cases from validation set to demonstrate the impact of different GRU configurations on prediction accuracy (Figure 4). We compare the GRU, ConvGRU(3), and ConvGRU(2+3) models across cases (a) to (d).
  • As shown in Case (a), the original GRU incorrectly predicts the answer to be "blue", likely confusing the color of the background with that of the text. Conversely, both the ConvGRU(3) and the ConvGRU(2+3) models accurately identify the correct answer as "white". This illustrates the capability of these enhanced models to capture key phrase pairs such as "color of" and "wording on" through 2-grams, and richer context through 3-grams like "the color of" and "of the wording," which effectively utilize the convolutional layers to discern the finer details of the question.
  • In Case (b), the GRU model gives a completely wrong prediction of "picture" as it relies solely on simplistic word-level features. The ConvGRU(3) model improves by predicting "sitting"—a rationally deduced yet contextually incorrect response, revealing that 3-gram features effectively capture a person’s state or action in a specific location. The ConvGRU(2+3), by incorporating both 2-gram and 3-gram features, correctly comprehends the query’s intent and the subject’s context, leading to an accurate "reading" prediction. This case exemplifies how combining different n-grams can enrich linguistic feature understanding and enhance alignment with image features for more accurate predictions.
  • In both the counting task and the yes/no task, as shown in Case (c) and Case (d), ConvGRU(2+3) showcases best performance. The ability to understand relational context and specific details allows it to correctly count "3" boys on the field and accurately confirm the presence of avocado on the sandwich with a "yes". These outcomes illustrate the effectiveness of integrating multi-gram features into GRUs.
Across all cases, the results consistently demonstrate that ConvGRU(2+3) achieves higher accuracy compared to ConvGRU(3) and the original GRU. This highlights the effectiveness of integrating multi-gram features in enhancing the GRU’s comprehension of textual information, thereby improving VQA accuracy without significantly increasing computational costs.
@@ 5 Conclusion
This paper explored the effectiveness of complex versus simpler models for capturing textual features in Visual Question Answering (VQA) tasks, specifically addressing whether complex sequential models are optimal for this purpose. Our experiments demonstrated that the ConvGRU model, designed to capture local text features, marginally outperformed the baseline GRU model, while advanced models like Self-Attention and Transformer Encoders showed reduced effectiveness. The primary advantage of ConvGRU lies in its ability to enhance performance with minimal computational cost, and it supports the importance of balancing model complexity with efficiency.
Nonetheless, our model is limited in scope, as it does not address other components of the VQA framework, such as the image encoder, where more advanced models beyond the Region-based Convolutional Neural Networks (R-CNN) could potentially improve results. Future research should focus on targeted improvements within both individual components and fusion methods to further optimize multi-modal understanding. Potential directions include exploring advanced neural architectures tailored for visual data and developing fusion techniques that can more effectively integrate visual and textual modalities to foster a deeper and more contextually relevant understanding in VQA tasks.